OpenStudy (anonymous):

approximate the relative minimum and/or maximum of f(x)=x^3-3x^2-12x

OpenStudy (anonymous):

sorry i wrot eit wrong

OpenStudy (anonymous):

approximate the relative minimum and/or maximum of f(x)=x^3-3x^2+2

OpenStudy (mimi_x3):

have you differentiated it?

OpenStudy (mimi_x3):

you differentiate it and set it to \(0\) for the max.

OpenStudy (mimi_x3):

\[f'(x) = 3x^2-6x\]

OpenStudy (mimi_x3):

then set it to \(0\)

OpenStudy (mimi_x3):

\[3x^2-6x=0\]

OpenStudy (mimi_x3):

i think that you're able to do it now :)

OpenStudy (anonymous):

what about the minimum?

OpenStudy (mimi_x3):

it's says max or min; you determine if its a max. or min..

OpenStudy (anonymous):

f(x)=x^3-3x^2-12x m=3x^2-6x-12 m'=6x-6 m'=0 x=1 therefore its min

OpenStudy (anonymous):

so there;s a minimum and maximum?

OpenStudy (mimi_x3):

wait you differentiate it twice to determine if its a max or min i think

OpenStudy (mimi_x3):

you solve for \(x\) first from \(f'(x)\) and to determine the nature i.e is the max/min you use the second derivative subbing in the \(x\) stationary point